OpenStudy (anonymous):

Check if the following infinite series is convergent or not (1/2)+((1/3)*(1+1/2))+((1/4)*(1+1/2+1/3))+.....

OpenStudy (amistre64):

can we create a summation rule for it?

OpenStudy (amistre64):

or distribute it out and make multiple summations ....

OpenStudy (anonymous):

we can write it as sigma ((1/(n+1))*(1+1/2+1/3+.......1/n))

OpenStudy (anonymous):

summation goes from 1 to infinity

OpenStudy (amistre64):

\[1(\frac{1}2{}+\frac{1}{3}+\frac{1}{4}+...)\] \[\frac12(\frac{1}{3}+\frac{1}{4}+\frac15+...)\] \[\frac14(\frac{1}{4}+\frac15+\frac{1}{6}+...)\] is what i think im seeing

OpenStudy (amistre64):

yours might be more apt for a ratio test tho

OpenStudy (anonymous):

yes but i am unable to write the (1+1/2+1/3+.....) term in short formula to apply that.

OpenStudy (amistre64):

\[a_n=\frac{1}{2^{n+1}}\sum_{i=0}^{n}\frac{1}{(i+1)}\]

OpenStudy (anonymous):

how did you write it?

OpenStudy (amistre64):

same as yours, but im thinking if an starts at 1, then 2^(n) might be more appropriate

OpenStudy (amistre64):

still not sure how well this will do for us .... the more basic idea is the distribution and collection of 'like' terms

OpenStudy (amistre64):

(1/2)+((1/3)*(1+1/2))+((1/4)*(1+1/2+1/3))+..... .... 1(1/2 + 1/3 + 1/4 + 1/5 + 1/6 + ...) 1/2(1/3 + 1/4 + 1/5 + 1/6 + ...) 1/4(1/4 + 1/5 + 1/6 + ...) 1/8(1/5 + 1/6 + ...) then the limit of a sum is the sum of the limits

OpenStudy (amistre64):

i dont see that any of the parts converges .... the first part is the harmonic and doesnt converge ,,, i dont see the others converging either but i cant prove it at the moment

OpenStudy (anonymous):

well actually this is a part of a bigger problem. We were supposed to check the convergence of the above series and another series just like above but with alternating terms which means : (1/2)-((1/3)*(1+1/2))+((1/4)*(1+1/2+1/3))-..... there are alternating signs with each term. This one i think can be done with liebnitz test.

OpenStudy (amistre64):

im not familiar with a liebnitz test so i cant verify that

OpenStudy (anonymous):

It is a test used to check the convergence of alternating infinite series. I spelled it wrong. It is "Leibnitz test".

OpenStudy (amistre64):

the first one posted is similar to the harmonic series, but which some values doubled, tripled etc so its larger by defualt and diverges. no idea about the alternating one :)

OpenStudy (anonymous):

right ,thanks